# Towards Immersive Mixed Reality Street Play

### v1

@@ 6.2.2. Ethical, Safety and Privacy Concerns in Public MR Play
Despite the positive social engagement observed during IMRSP experiences, our findings indicate significant ethical, safety, and privacy concerns (Katins et al., 2024; Millard et al., 2024) that must be addressed to facilitate broader social acceptance (Eghbali et al., 2019). A prominent issue is the power asymmetry between MR headset wearers and non-participating bystanders. MR users have privileged access to hidden digital information through outward-facing sensors, such as cameras and depth scanners, potentially recording bystanders without explicit consent or awareness. This dynamic creates a sense of vulnerability among non-users, who might perceive MR technology as intrusive or threatening their privacy—an effect reminiscent of earlier public reactions to technologies like Google Glass, where similar privacy concerns led to significant public backlash (Pelikan et al., 2024; McGrane, 1994).
[…]
@@
Physical safety represents another critical concern identified in our research. Although no incidents occurred during our field studies, the immersive, fast-paced nature of IMRSP raises genuine risks of accidents or collisions. Players engrossed in virtual gameplay might inadvertently overlook physical hazards, such as vehicles, street furniture, or pedestrians. Online observers particularly highlighted these potential dangers, reflecting valid concerns about player distraction in dynamic public environments. Additionally, social safety—how comfortable and secure bystanders feel around MR participants—must be carefully considered. Sudden or unpredictable movements by MR players could alarm or unsettle individuals who are unfamiliar with the activity, creating unintended social tension.
@@
Furthermore, IMRSP explicitly challenges established social norms, particularly Goffman’s concept of ”civil inattention,” where individuals typically ignore each other in public settings to maintain social order (Goffman, 2008). By introducing visible, unconventional behaviors into shared public spaces, MR street play inherently disrupts these existing norms, compelling bystanders to acknowledge and respond to unusual public performances. While some individuals might appreciate these playful disruptions as novel or refreshing, others could interpret them negatively, perceiving such activities as intrusive or socially inappropriate. This tension could potentially lead to social friction or resistance against MR enthusiasts, highlighting the need for proactive design and policy considerations to manage public interactions sensitively.
@@
Addressing these ethical and social challenges requires careful attention to transparency and consent regarding MR technology usage, ensuring clear communication about data capture and privacy implications. Emphasizing group-oriented rather than solitary gameplay may reduce perceptions of awkwardness or intrusiveness. Moreover, urban spatial planning that clearly defines appropriate public spaces or dedicated zones for MR activities could enhance both physical and social safety. Proactive measures addressing privacy, safety, and public etiquette concerns will be essential for the successful integration and acceptance of IMRSP in everyday urban life, preventing potential social backlash and fostering a harmonious coexistence between MR players and the broader public.
6.3. Design Recommendations for Socially-Acceptable IMRSP
@@
Based on our findings and the above discussion, we distill several key design suggestions for future IMRSP development.
6.3.1. Design for Contextual Appropriateness
Our findings suggest that MR street play fosters collocated interactions and public performance, which implies that MR design should enhance the social aspect of play. Designing MR applications with features that facilitate teamwork, shared challenges, and visible achievements can encourage group engagement and reduce the potential for social discomfort, particularly in settings with bystanders. Moreover, allowing for adaptive performative levels in embodied interaction, might encourage users to transition between subtle and more performative engagements depending on their comfort levels and social settings. Given that MR street play often turns the public space into a stage, designing for spectator experience, or public display and audience engagement can enhance the shared experience for both players and bystanders.
6.3.2. Harmonizing with Social Norms with Explainable Social Affordance
@@
As MR street play disrupts conventional social norms, it creates a disconnect between players’ social and virtual identities, potentially causing social awkwardness. There’s a need to design MR street play that promotes social etiquette and manages instances of ”civil inattention” (Goffman, 2008) in public settings. This can be achieved by incorporating narrative elements, role design, and physical props that signal ongoing MR activities to bystanders, thus reducing confusion and promoting mutual awareness. For example, in our MOFA game, we emphasized the use of physical props, such as wands, to rationalize players’ behavior. Given the narrative roles of wizards and witches, these props bridge the gap by offering a tangible connection to the player’s in-game persona, enhancing player immersion while alleviating bystanders’ concerns. These social affordances also help lower the learning curve for players. Furthermore, incorporating additional elements like music, costumes, and visual cues can enrich the gaming experience and enable clearer communication of player behaviors to the public. Furthermore, designers can help users navigate social boundaries by providing guidelines on appropriate behaviors within different social and cultural contexts. This might include features that notify users when they enter culturally sensitive areas or crowded spaces where MR play could potentially cause social friction.
6.3.3. Guiding Users to Appropriate Locations
@@
6.4. Limitations
This paper represents an initial exploration of IMRSP through our game probe MOFA, which we selected for its straightforward implementation using current XR device capabilities. While this approach yielded valuable insights, we acknowledge several limitations.
Limited Game Play
[…]
Limited Physical Interaction
Due to the game design and safety considerations, all three games in MOFA have no physical contact, which limited our exploration of potential physical interaction dynamics in MR street play. Physical contact could introduce richer, more complex social interactions that might enhance or complicate MR experiences in public. Future studies might explore MR street play scenarios that permit safe, controlled physical interactions to better understand how touch and close proximity impact social bonding, comfort, and behavioral dynamics in public MR environments.
Limited Environmental Interaction
The environmental interaction in our study was limited, as the MOFA games were designed primarily to encourage social interaction among players rather than interaction with the surrounding physical environment. While players could move freely within certain spaces, the games did not incorporate elements of the physical environment, such as urban infrastructure, landmarks, or other spatial features, which could enhance immersion and contextual relevance. Future studies could explore MR street play designs that integrate environmental interactions—such as using real-world objects as interactive components or adapting gameplay based on specific environmental characteristics.
6.5. Future Research Agenda
For future research on IMRSP, we propose the following agenda:
@@ 6.5. Future Research Agenda
[…]
Future researchers can explore more quantitative research on social acceptance factors blocking the widespread adoption of IMRSP. They can conduct A/B testing to understand why people initially reject this game online but accept it after experiencing it in person. They can even invite those who reject the game online to try it onsite, allowing them to study their reactions and long-term retention behaviors. This research would help identify the key barriers preventing this type of game from becoming successful.
@@
As this field remains largely uncharted, we encourage researchers to investigate its various aspects and potential. This includes exploring different game mechanics, interaction modalities, and social dynamics that are unique to IMRSP.
@@
These guidelines should address issues such as player privacy, physical safety in public spaces, and potential psychological impacts of immersive experiences in urban environments.
Develop protocols for inviting bystanders to join IMRSP
What permissions are necessary? How can we protect privacy? Which invitation methods are most effective? This research could involve testing various invitation strategies and analyzing their effectiveness and social acceptability.
Facilitate intent inclusivity in public spaces for spectatorship
As described in (Hu et al., 2024), this includes understanding the Intent Differentiation of HMD wearers and non-HMD wearers. For HMD wearers, this may include unconcerned players, privacy-conscious players, acceptable non-sharers, eager sharers, and casual inviters. For non-HMD wearers, this may include enthusiastic recruiters, indifferent passersby, tolerant observers, disturbed onlookers, curious spectators, included spectators, and those ready to join with handheld AR or head-mounted displays. Studying these different intents could lead to more inclusive and socially aware IMRSP experiences.
Beyond wands, this could involve designing and testing various physical props, visual cues, or auditory signals that help convey players’ actions and intentions to bystanders.
Explore the relationship between Live Action Role Play (LARP) and IMRSP
How can mixed reality enhance LARP (Johansson et al., 2024) experiences? This research could involve adapting existing LARP scenarios to IMRSP contexts and analyzing the benefits and challenges of this integration.
Investigate ways to incorporate more environmental interactions, such as with buildings, streets, and landscapes
This could include developing techniques for real-time environment mapping and creating dynamic game elements that respond to the urban landscape.
Theoretical exploration of IMRSP’s relationship to Ludology, Urban Games, Playful City, Urbanism, Persuasive Games, Mixed Reality Games, and Exertion Games
For future research on IMRSP, we suggest a theoretical exploration of its relationships with various fields. This includes investigating how IMRSP fits within existing game theories and urban play concepts, examining its potential to transform urban spaces into interactive playgrounds, exploring its use for social change or education in urban environments, and analyzing its intersection with immersive technologies and physical activity-based gaming.
@@ 7. Conclusion
[…]
Based on our first in-the-wild empirical studies of IMRSP using our open-source exploratory game probe series MOFA, we found that MOFA demonstrates IMRSP’s potential as a new paradigm of highly social and engaging exertion-based play in public spaces, as verified through observations of social phenomena and dynamics. By comparing online reviews with onsite gameplay experiences, we identified key challenges in social acceptance and environmental factors that influence adoption. These insights inform our suggested next steps and research agenda for IMRSP. We concluded that despite its novelty and highly engaging social nature, the widespread acceptance of IMRSP in public spaces may take longer than anticipated due to the inherent power imbalance created by MR headsets.

### v2

@@ 6.2.2. Ethical, Safety, and Privacy Concerns in Public MR Play
Despite the positive social engagement observed during IMRSP experiences, our findings indicate significant ethical, safety, and privacy concerns (Katins et al., 2024; Millard et al., 2024) that must be addressed to facilitate broader social acceptance (Eghbali et al., 2019). A prominent issue is the asymmetric power between MR headset wearers and non-participating bystanders. MR users have privileged access to hidden digital information through outward-facing sensors, such as cameras and depth scanners, potentially recording bystanders without explicit consent or awareness. This dynamic creates a sense of vulnerability among non-users, who might perceive MR technology as intrusive or threatening their privacy—an effect reminiscent of earlier public reactions to technologies like Google Glass, where similar privacy concerns led to significant public backlash (Bajorunaite et al., 2024). Recent research also underscores the necessity for privacy-enhancing mechanisms that account for bystanders’ varying preferences for awareness and consent in everyday AR scenarios (O’Hagan et al., 2022).
@@
Physical safety represents another critical concern identified in our research. Although no incidents occurred during our field studies, the immersive, fast-paced nature of IMRSP raises genuine risks of accidents or collisions. Players engrossed in virtual gameplay might inadvertently overlook physical hazards, such as vehicles, street furniture, or pedestrians. Online observers particularly highlighted these potential dangers, reflecting valid concerns about player distraction in dynamic public environments. Additionally, social safety—how comfortable and secure bystanders feel around MR participants—must be carefully considered. Sudden or unpredictable movements by MR players could alarm or unsettle individuals who are unfamiliar with the activity, creating unintended social tension. As social MR systems grow more pervasive, the adoption and effectiveness of in-situ safety tools to mitigate harassment, discomfort, or unwanted encounters become increasingly crucial for ensuring both player and bystander well-being in shared environments (Weerasinghe et al., 2025).
@@
Furthermore, IMRSP explicitly challenges established social norms, particularly Goffman’s concept of “civil inattention,” where individuals typically ignore each other in public settings to maintain social order (Goffman, 1966). By introducing visible, unconventional behaviors into shared public spaces, MR street play inherently disrupts these existing norms, compelling bystanders to acknowledge and respond to unusual public performances. While some individuals might appreciate these playful disruptions as novel or refreshing, others could interpret them negatively, perceiving such activities as intrusive or socially inappropriate. This tension could potentially lead to social friction or resistance against MR enthusiasts, highlighting the need for proactive design and policy considerations to manage public interactions sensitively. Additional scholarship highlights that the widespread adoption of everyday AR will introduce significant ethical challenges—such as manipulation and information disorder—necessitating clear guidelines for who can mediate reality and for what purpose (O’Hagan et al., 2023). Building on this, recent work argues that everyday, pervasive AR demands a new societal consensus on “perceptual rights” to protect both users and bystanders from potential misuse and abuse (O’Hagan et al., 2024).
@@
Addressing the social and ethical challenges in IMRSP requires careful attention to transparency and consent regarding MR technology usage, ensuring clear communication about data capture and privacy implications. Emphasizing group-oriented rather than solitary gameplay may reduce perceptions of awkwardness or intrusiveness (Deterding, 2018). Moreover, urban spatial planning that clearly defines appropriate public spaces or dedicated zones for MR activities could enhance both physical and social safety. Proactive measures addressing privacy, safety, and public etiquette concerns will be essential for the successful integration and acceptance of IMRSP in everyday urban life, preventing potential social backlash and fostering a harmonious coexistence between MR players and the broader public.
6.3. Methodological Implications for MR Futuring
In this paper, we use today’s MR technology to enact MR’s own futures. Our methodology employs a fully-functional high-fidelity research probe that instantiates a plausible next-step MR experience in situ. This enables observation of real social consequences—bystander dynamics, on-the-fly norm negotiation, and emergent coordination—rather than relying on hypothetical narration or studio vignettes. Methodologically, we blend research-through-design (Zimmerman et al., 2007) with immersive speculative enactment (Simeone et al., 2022). However, our approach differs from prior immersive speculative enactment works that typically use VR to depict futures from within a headset. Instead, we use today’s MR itself to probe MR’s own future, deployed in public space with real passersby. This yields ecologically valid evidence about how people move, signal, and coordinate around MR systems—evidence that is hard to surface in purely fictional or studio-bound enactments.
This approach is effective because, although MR hardware will plausibly improve (e.g., smaller form factors, wider FOVs, better tracking), the core social-cognitive frictions that emerge in public, multi-party settings are likely to persist. MR technology-mediated experiences obscure what others are attending to, hindering theory of mind (McKenna et al., 2023) and making norms less visually legible to bystanders who cannot see what the wearer sees (Von Terzi et al., 2021; Desnoyers-Stewart et al., 2025). Passthrough and layered visualizations further complicate mutual understanding in public places (Bailenson et al., 2024). Additionally, MR renders reality programmable: MR designers effectively codify “strange new rules” (Rao and Protocolized, 2025)—what we call “reality protocols” (Hu, 2025)—that determine what counts as visible, actionable, or socially appropriate within an app. Because these protocols vary by app, they can drive divergent behaviors in the same physical place, further hindering theory of mind even between MR wearers. This limitation will likely persist even as hardware improves. The embodied signatures we observe—unshared referents, ambiguous gestures, and micro-breaches of street etiquette—are system-agnostic. They emerge from asymmetries of technology-mediated perception and accountability rather than from any single device’s limitations. Consequently, findings from today’s deployments speak to the long arc of MR sociality: as people encounter device-specific, app-specific rules that aren’t universally visible, they will continue to negotiate trust, accountability, and shared context in public.
Fortunately, thanks to MR’s programmability, many of tomorrow’s rules can be implemented today: we can encode prospective interaction logics, object behaviors, and social norms on current hardware; deploy them in real environments; and study how trust, accountability, and shared context are actually negotiated. In essence, MR offers a prefigurative testbed: by programming futures into the present, we can derive empirically grounded implications for those futures, echoing William Gibson’s famous insight:
The future is already here—it’s just not very evenly distributed.
Methodologically, this paper suggests three commitments for MR futuring. First, favor enactment over narration: stage near-future MR experiences in public settings where bystander interpretation, coordination, and etiquette are genuinely at stake. Second, treat MR prototypes as norm engines: vary the reality protocol (what is visible/doable/appropriate) and observe how rule changes reconfigure gesture legibility, referent alignment, and accountability. Third, analyze what persists: focus on social frictions rooted in asymmetric perception and theory-of-mind constraints, which are likely to endure across device generations.
In sum, by implementing speculative rule sets now and examining their social consequences in real-world settings, our method provides a concrete pathway for researchers and designers to co-produce MR futures with the public. This approach goes beyond merely imagining possibilities—it allows us to observe how these futures actually unfold in practice. We anticipate this methodology will inspire additional work that engages users in prospective MR frameworks, thereby advancing both design practice and our empirical understanding of emerging social norms.
6.4. Design Recommendations for Socially Acceptable IMRSP
@@
Based on our game development experience, findings, and the above discussion, we distill several key design suggestions for future IMRSP development.
6.4.1. Designing for Contextual Awareness
Given its co-located interactions in public spaces, we propose that IMRSP designs should explicitly consider contextual awareness (Grubert et al., 2017) across multiple dimensions. First, designers should critically evaluate the social norms and cultural acceptability of proposed embodied interactions within specific public contexts. This includes designing game themes appropriate for these contexts. Second, embracing the multiplayer aspect of IMRSP, mechanics that encourage collaborative engagement not only enhance the play experience but also mitigate social awkwardness through collective participation. Third, designs should incorporate adaptable interaction thresholds that allow users to modulate their performative engagement based on environmental factors and personal comfort levels. Fourth, since IMRSP transforms public spaces into performative arenas, designers should intentionally craft spectator experiences (Reeves et al., 2005) that facilitate understanding and potential participation. Finally, we recommend that designers systematically analyze contextual parameters—including spatial characteristics, cultural specificity, and social dynamics—before designing IMRSP applications in public settings.
6.4.2. Harmonizing with Social Norms through Explainable Social Affordance
@@
As MR street play disrupts conventional social norms, it creates a disconnect between players’ social and virtual identities, potentially causing social awkwardness. There is a need to design MR street play that promotes social etiquette and manages instances of civil inattention (Goffman, 1966) in public settings. This can be achieved by incorporating narrative elements, role design, and physical props that signal ongoing MR activities to bystanders, thus reducing confusion and promoting mutual awareness. For example, in our MOFA game, we emphasized the use of physical props, such as wands, to rationalize players’ behavior. Given the narrative roles of wizards and witches, these props bridge the gap by offering a tangible connection to the player’s in-game persona, enhancing player immersion while alleviating bystanders’ confusion. These social affordances (Isbister et al., 2018; Wu et al., 2020) also help lower the learning curve for players. Furthermore, incorporating additional elements like music, costumes, and visual cues can enrich the gaming experience and enable clearer communication of player behaviors to the public. Finally, designers can help users navigate social boundaries by providing guidelines on appropriate behaviors within different social and cultural contexts. This might include features that notify users when they enter culturally sensitive areas or crowded spaces where MR play could potentially cause social friction.
6.4.3. Designing Engaging Movement while Maintaining Safety
IMRSP naturally engages users’ movement as its interaction mechanism. We recommend four design considerations to ensure both engaging gameplay and public safety. First, maintain a largely unobstructed, transparent real-world view in the MR interface to help players remain aware of their physical surroundings, thereby reducing collision risks when moving in public spaces. For example, in MOFA, augmented overlays are intentionally minimal to prevent users from losing their sense of reality: most of the display area in MR should remain free of UI elements and augmented content. Second, design interaction mechanics that limit potentially dangerous movements—for example, MOFA’s magic shield system provides clear attack targets without requiring physical contact between players. Third, incorporate visual feedback systems that substitute for physical impacts or UI elements. For example, the shield-breaking animations signify instant feedback of attack, avoiding the need to overlay traditional health bars over opponents. This creates visceral feedback without requiring players to frequently check UI elements for status updates, which would distract attention and potentially cause fatigue. Finally, include spatial boundary indicators that dynamically adjust based on surrounding pedestrian density, helping players maintain appropriate play zones in varying public conditions. These design principles help balance the inherently movement-based nature of IMRSP with the safety considerations necessary for deployment in shared public environments.
6.4.4. Guiding Users to Appropriate Gameplay Locations
@@
6.5. Limitations and Future Work
This paper presents an initial exploration of IMRSP through our game probe MOFA, demonstrating the capabilities of current affordable XR devices to support MR street play in the wild. While this approach yielded valuable insights, we acknowledge several limitations and propose a research agenda for future work on IMRSP.
6.5.1. In‑the‑Wild Scenarios
In this study, we examined only five scenarios where IMRSP was deployed, which may not fully represent everyday “in-the-wild” conditions. For instance, in exhibition contexts such as expos and conferences, participants were predisposed to encounter novel technologies, creating an atmosphere of heightened tolerance for unconventional embodied interactions. Similarly, campus settings, where physical activities are commonplace (e.g., frisbee games), may have provided a more accommodating social context than truly public spaces. These settings likely attracted participants with greater openness toward experimental gameplay, potentially mitigating social acceptability concerns that might arise in genuinely public environments. This contextual bias may partially explain the divergent perspectives between in-person and online commenters, where online respondents emphasized the potentially disruptive aspects of the game, while in-person participants primarily focused on positive elements, experiencing minimal social awkwardness as others actively engaged in the experience. As this research represents an initial exploration of IMRSP, future studies should extend to more ecologically valid scenarios similar to how Pokémon GO operates across diverse public environments, to better understand the social dynamics and acceptance of such technologies in authentic everyday settings. Future work could look into the following directions:
@@ 6.5.1. In‑the‑Wild Scenarios
Future researchers can explore longitudinal research on social acceptance factors blocking the widespread adoption of IMRSP. It would be meaningful to understand why people initially reject this game online but accept it after experiencing it in person. They can even invite those who reject the game online to try it on-site, allowing them to study user reactions and long-term retention behaviors. This research would help identify the key barriers preventing this type of game from becoming successful.
Facilitate intent inclusivity in public spaces for spectatorship
As described in (Hu et al., 2024a), facilitating intent inclusivity includes understanding the intent differentiation of HMD wearers and non-HMD wearers. For HMD wearers, this may include unconcerned players, privacy-conscious players, acceptable non-sharers, eager sharers, and casual inviters. For non-HMD wearers, this may include enthusiastic recruiters, indifferent passersby, tolerant observers, disturbed onlookers, curious spectators, included spectators, and those ready to join with handheld AR or head-mounted displays. Studying these different intents could lead to more inclusive and socially aware IMRSP experiences.
6.5.2. Gameplay
Our study’s MR street play experiences were based on a single game probe—MOFA, which features three games within a magic-themed context. While this theme provided a shared framework for players to quickly understand gameplay mechanics, it may not fully capture the diverse range of potential MR street play scenarios. Although we tested only one magical fantasy theme, we believe this approach can be applied more broadly. Future studies should expand beyond magic and explore games with different themes, interaction styles, and levels of physical engagement to better understand the full spectrum of MR street play possibilities and their effects on social dynamics. We propose the following research agenda:
@@ 6.5.2. Gameplay
This field remains largely uncharted, so we encourage researchers to investigate its various aspects and potential. This includes exploring different game mechanics, interaction modalities, and social dynamics unique to IMRSP.
Similar to the wands used in MOFA, future work should design and test various physical props, visual cues, and auditory signals that help convey players’ actions and intentions to bystanders.
6.5.3. Physical Interaction
Due to the game design and safety considerations, all three games in MOFA required no physical contact, which limited our exploration of potential physical interaction dynamics in MR street play. We acknowledge the potential of physical contact in introducing richer, more complex social interactions, which might enhance or complicate MR experiences in public. Future studies might explore MR street play scenarios that permit safe, controlled physical interactions to better understand how touch and close proximity impact social bonding, comfort, and behavioral dynamics in public MR environments. In response to this limitation, future work should:
@@ 6.5.3. Physical Interaction
These guidelines should address issues such as player privacy, physical safety in public spaces, and the potential psychological impacts of immersive experiences in urban environments. In particular, developing protocols for inviting bystanders to join IMRSP is essential. What permissions are necessary? How can we protect privacy? Which invitation methods are most effective? This research could involve testing various connection strategies and analyzing their effectiveness and social acceptability.
6.5.4. Environmental Interaction
The environmental interaction in our study was limited, as the MOFA games were designed primarily to encourage social interaction among players rather than interaction with the surrounding physical environment. While players could move freely within certain spaces, the games did not incorporate elements of the physical environment, such as urban infrastructure, landmarks, or other spatial features, which could enhance immersion and contextual relevance. Future studies could explore MR street play designs that integrate environmental interactions, such as using real-world objects as interactive components or adapting gameplay based on specific environmental characteristics.
Theoretical exploration of IMRSP’s relationship to adjacent forms of play and games
For future research on IMRSP, we suggest a theoretical exploration of its relationships with various fields, such as social games (Gonçalves et al., 2023), urban games (De Luca and Bertolo, 2012; Woolley and Lowe, 2013), alternate reality games (Kim et al., 2008), playful city (Nijholt, 2017a), playful wearables (Buruk et al., 2023), pervasive games (Ahn et al., 2016), location-based games (Laato et al., 2019; Sutko and De Souza E Silva, 2011), mixed reality games (Bonsignore et al., 2012; Millard et al., 2024), live action role play (LARP) (Johansson et al., 2024; Márquez Segura et al., 2018), sport games (Mueller et al., 2011), and exertion games (Mueller et al., 2008, 2016). This includes investigating how IMRSP aligns with existing game theories and urban play concepts, examining its potential to transform urban spaces into interactive playgrounds, exploring its use for social change or education in urban environments, and analyzing its intersection with immersive technologies and physical activity-based gaming.
@@ 7. Conclusion
This work presents our in-the-wild empirical studies of immersive mixed reality street play (IMRSP) using our open-source exploratory game probe series MOFA. Our findings demonstrate IMRSP’s potential as a new paradigm for highly social and engaging exertion-based play in public spaces, supported by observations of rich social phenomena and dynamics. By comparing online reviews with on-site gameplay experiences, we identified key challenges in social acceptance and environmental factors affecting adoption. These insights inform our suggested next steps and research agenda for IMRSP. We conclude that, despite its engaging social nature, widespread acceptance of IMRSP may take longer than anticipated due to social, ethical, safety, and privacy concerns.
Acknowledgements.
This research was funded by Holo Interactive. We thank Yuchen Zhang and Sizheng Hao from Holo Interactive for developing the MOFA system. We extend our gratitude to Professor Steve Benford and Professor Gisela Reyes-Cruz from Mixed Reality Lab at University of Nottingham for their valuable advice and suggestions. We especially thank Eva Yutzu Chen for assisting with user testing. We are grateful to students from the School of Design and Innovation at China Academy of Art, IDM and ITP students from New York University, and 706 Youth Community in Dali, China for supporting our user tests. Additionally, we appreciate the participants from SIGGRAPH, CHI, CHI Play, and UbiComp/ISWC onsite exhibitions for their valuable feedback that helped improve the system.